# Multimodal foundation world models for generalist embodied agents

### v1

@@ Appendix A Extended Related Work
Foundation models for actions. Few cases of foundation models for embodied domains have been developed. Notable mentions are GATO [44], a large-scale behavior cloning agent, trained on 604 tasks. VPT [2] a large-scale model trained on Minecraft data, using human-expert labeled trajectories. The model learns strong behavioral priors by behavior cloning which can be fine-tuned using RL. STEVE-1 [34] connects VPT’s behavioral prior with the MineCLIP model representation [15], using the unCLIP approach [43]. RT-X [12] are large-scale trasnformer models trained on expert robotics dataset, sharing a common action space (end-effector pose) across different embodiments.
@@ Appendix B Implementation details
The connector network uses the same hyperparameters and architecture as the sequential dynamics of the world model. The aligner network employs a small U-Net, with a bottleneck that is half the size of the embedding representation.
@@ Appendix C Tasks
We present the list of tasks employed, along with the language prompts used for specifying the task, in Table 3. We introduce a new embodied environment, the ‘Stickman’, which serves as a humanoid robot (compared to the one present in the dm_control suite) that is simpler to control, thanks to a reduced number of joints. Its addition allows studying behaviors that involve upper body movements, rather than focusing on lower body motions. For the newly introduced tasks, the goal can be easily inferred by reading the task’s name or its prompt. For the ‘flipping’ tasks, we consider flips both in forward direction and backward direction, as the VLM struggles to distinguish directions.
@@ Appendix D Experiments settings
Baselines. In order to implement performant offline RL baselines we adopt the findings of [49] and [7], adopting larger deeper networks and layer normalization.
Inputs are 64x64x3 RGB images. We use a frame stack of 3. The encoder architecture is adapted from the DrQ-v2 encoder [59]. We did find augmentations on the images, e.g. random shifts, to hurt performance.
Offline RL. For each task, training model-free agents (IQL, TD3, TD3+BC) requires re-training the full agent (visual encoder, actor, critic) on the entire dataset, from scratch, while training model-based agents (GenRL) requires training the model once for each domain and then training an actor-critic for each task. Moreover, for training the actor-critic in GenRL, we only use 50k gradient steps, as the policy converges significantly faster than for the other methods.
Compute resources. We use a cluster of V100 with 16GB of VRAM for all our experiments. To enable efficient training, image and video-CLIP embeddings are computed in advance and stored with the datasets. Training the MFWM for 500k gradient steps takes $\sim$ 5 days. After pre-training the MFWM, training the actor-critic for a prompt for 50k gradient steps takes less than 5 hours. In data-free mode, it takes less than 3 hours. In both cases, convergence normally arrives after 10k gradient steps, but we keep training. Model-free baselines take around 7 hours to train for 500k gradient steps. normally arrive
@@
Appendix E Additional results
@@
Figure 7: Temporal alignment ablation. We analyze the impact of temporal alignment in our proposed RL objective for matching sequential targets. Results averaged over 3 seeds.
@@
Figure 8: Multi-task generalization detailed results. Results averaged over 5 seeds.
@@
Figure 9: Data-free RL detailed results. Results averaged over 5 seeds.
@@
Figure 10: Training data distribution detailed results. Results averaged over 3 seeds.

### v2

@@ Appendix A Extended Related Work
Foundation models for actions. Few cases of foundation models for embodied domains have been developed. Notable mentions are GATO [47], a large-scale behavior cloning agent, trained on 604 tasks. VPT [2] a large-scale model trained on Minecraft data, using human-expert labeled trajectories. The model learns strong behavioral priors by behavior cloning which can be fine-tuned using RL. STEVE-1 [35] connects VPT’s behavioral prior with the MineCLIP model representation [15], using the unCLIP approach [46]. RT-X [12] are large-scale transformer models trained on expert robotics dataset, sharing a common action space (end-effector pose) across different embodiments.
@@ Appendix B Implementation details
The connector network uses the same hyperparameters and architecture as the sequential dynamics of the world model. The aligner network employs a small U-Net, with a bottleneck that is half the size of the embedding representation. The embedding representation is 768-dimensional. Further details can be found in our accompanying code implementation.
@@ Appendix C Tasks
Stickman environment. The Stickman environment is based on the Walker environment from the dm_control suite. We designed the Stickman environment to explore tasks that require upper body limbs (e.g. boxing, doing a handstand) without the complexity of training a humanoid (which requires a significantly larger amount of data to be solved [63]). The number of joints is increased by 4: 2 joints per arm, one is for the shoulder, the other for the elbow. The total number of joints is 10. The action space is normalized to be in [-1,1] as all dm_control tasks. The robot also presents a head, to resemble a humanoid.
Prompts and scores. We present the list of tasks employed, along with the language prompts used for specifying the task, in Table 3. For the newly introduced tasks, the goal can be easily inferred by reading the task’s name or its prompt. For the ‘flipping’ tasks, we consider flips both in the forward direction and backward direction, as the VLM struggles to distinguish directions. The reward functions used to evaluate the agent’s score can be found in our open-source code.
@@ Appendix D Experiments settings
Baselines. In order to implement performant model-free offline RL baselines we adopt the findings of [52] and [7], adopting larger deeper networks and layer normalization. Inputs are 64x64x3 RGB images. We use a frame stack of 3. The encoder architecture is adapted from the DrQ-v2 encoder [63]. We did find augmentations on the images, e.g. random shifts, to hurt performance.
The WM-CLIP baseline learns a “reversed connector" from the world model representation to the VLM representation (GenRL does the opposite). The “reversed connector", given the latent state corresponding to a certain observation, predicts the corresponding embedding. Formally:
$\begin{split}\textrm{Reversed connector:}\quad\hat{e}^{(v)}_{t}=f_{\psi}(s_{t},h_{t})\end{split}\quad\quad\begin{split}\mathcal{L_{\textrm{rev\_conn}}}=\|e^{(v)}-\hat{e}^{(v)}_{t}\|^{2}_{2}\end{split}$
After training the reversed connector, visual embeddings can be inferred from latent states. For policy learning, rewards are computed using the cosine similarity between embeddings inferred from imagined latent states and the prompts’ embedding.
The reversed connector is implemented as a 4-layer MLP, with hidden size 1024. For fair comparison, we adopt the same world model for WM-CLIP and GenRL. For WM-CLIP we pre-train the additional reversed connector, while for GenRL the connector and aligner.
Offline RL. For each task, training model-free agents (IQL, TD3, TD3+BC) requires re-training the full agent (visual encoder, actor, critic) on the entire dataset, from scratch, while training model-based agents (GenRL, WM-CLIP) requires training the model once for each domain and then training an actor-critic for each task. Moreover, for training the actor-critic in GenRL, we only use 50k gradient steps, as the policy converges significantly faster than for the other methods.
@@
Compute resources. We use a cluster of V100 with 16GB of VRAM for all our experiments. To enable efficient training, image and video-CLIP embeddings are computed in advance and stored with the datasets. Training the MFWM for 500k gradient steps takes $\sim$ 5 days. After pre-training the MFWM, training the actor-critic for a prompt for 50k gradient steps takes less than 5 hours. In data-free mode, it takes less than 3 hours. In both cases, convergence normally arrives after 10k gradient steps, but we keep training. Model-free baselines take around 7 hours to train for 500k gradient steps.
On a single GPU, model-free RL is faster to train for a small number of runs. GenRL starts becoming advantageous when using the world model for training for more than 60 runs (which is often the case, considering the number of runs = N seeds x M tasks per domain). When adopting the data-free policy learning strategy, GenRL doesn’t rely on the dataset at all. This halves the time required for training, as there are no data transfers between the CPU (where the dataset is normally loaded) and the GPU for training.
Appendix E Temporal alignment ablation
In Figure 8, we investigate the impact of our best-matching trajectory strategy for temporally aligning imaginary latent states to target states for GenRL’s reward computation.
@@ Appendix E Temporal alignment ablation
Figure 8: Temporal alignment ablation. We analyze the impact of temporal alignment in our proposed RL objective for matching sequential targets. Results averaged over 10 seeds.
Appendix F Aligner model ablation study
To establish the importance of the aligner network, in Table 5 we report the results of additional ablations:
  • GenRL - no aligner: this is an ablation of GenRL where the the language prompt’s embedding is directly fed into the connector, rather than processing it first with the aligner;
  • TD3-V and WM-CLIP-V + aligner: for these ablations, we first process the language prompt’s embedding using GenRL’s pre-trained aligner. Then, we use it to compute the cosine similarity for the reward function, as for the original baselines.
Table 5: Aligner ablation study.
GenRL - no aligner GenRL WM-CLIP-V WM-CLIP-V + aligner TD3-V TD3-V + aligner
cheetah 0.32 ± 0.02 0.93 ± 0.01 0.82 ± 0.17 0.84 ± 0.02 0.31 ± 0.09 0.77 ± 0.04
stickman 0.09 ± 0.01 0.66 ± 0.01 0.54 ± 0.03 0.51 ± 0.03 0.38 ± 0.02 0.38 ± 0.02
walker 0.19 ± 0.01 0.75 ± 0.01 0.70 ± 0.02 0.74 ± 0.01 0.56 ± 0.04 0.48 ± 0.04
kitchen 0.25 ± 0.00 0.76 ± 0.08 0.71 ± 0.14 0.84 ± 0.09 0.32 ± 0.16 0.27 ± 0.10
quadruped 0.17 ± 0.02 0.91 ± 0.02 0.81 ± 0.04 0.76 ± 0.05 0.32 ± 0.04 0.33 ± 0.04
overall 0.17 ± 0.00 0.76 ± 0.01 0.67 ± 0.03 0.68 ± 0.02 0.40 ± 0.02 0.42 ± 0.02
We can observe that: i) the aligner mechanism is crucial in GenRL’s functioning. ii) processing the language embedding in the reward function of the WM-CLIP-V and TD3-V baselines changes performance on some tasks (performance per domain varies). However, using the aligner provides no advantage overall.
We believe the aligner is very important in GenRL because its output, the processed language embedding, is fed to another network, the connector. If the language embeddings were not processed by the aligner, they would have been too different from the embeddings used to train the connector, which are the visual embeddings.
Instead, for the baselines, we process the language embedding with the aligner and then use it to compute a similarity score with the visual embeddings. This overall renders very similar performance to no aligner processing, hinting that the aligner network doesn’t improve the cosine similarity signal. At the same time, this also suggests that the aligner network doesn’t hurt the generality of the VLM’s embeddings, as the cosine similarity after processing the embedding provides a similarly useful signal as before processing.
Appendix G Comparison with LIV
We also tested all baselines with the LIV’s representation [40] in the Kitchen tasks. We used LIV’s open-source code to download and instantiate the model. Note that the available model is the general pre-trained model, not the one fine-tuned for the Kitchen environment.
LIV’s results confirm the original paper’s claims (see Appendix G4) that the representation does not work well for vision-language rewards without fine-tuning on domain-specific vision-language pairs (which are unavailable in our settings, as we use no language annotations).
Table 6: Comparison with LIV [40] on kitchen tasks.
IQL + LIV TD3+BC + LIV TD3 + LIV WM-CLIP + LIV GenRL
kitchen microwave 0.03 ± 0.03 0.0 ± 0.0 0.2 ± 0.16 0.0 ± 0.0 0.97 ± 0.02
kitchen light 0.53 ± 0.24 0.05 ± 0.04 0.0 ± 0.0 0.0 ± 0.0 0.46 ± 0.09
kitchen burner 0.2 ± 0.08 0.0 ± 0.0 0.0 ± 0.0 0.0 ± 0.0 0.62 ± 0.07
kitchen slide 0.03 ± 0.03 0.03 ± 0.03 0.0 ± 0.0 0.67 ± 0.27 1.0 ± 0.0
overall 0.20 ± 0.11 0.02 ± 0.02 0.05 ± 0.07 0.17 ± 0.12 0.76 ± 0.08
Appendix H Extended Discussion on Data-free Policy Learning
Figure 9: Representing how the different agents process the data for policy learning. By removing data dependencies on actions (on-policy learning in the model’s imagination), rewards (computed using only the prompt and latent states, using Eq. 3), and observations (by sampling latent states within the model), GenRL agents can be adapted for new tasks in a data-free fashion.
Offline RL methods (Fig. 9a) learn from offline datasets of observations, actions and rewards.
Offline RL methods (Fig. 9b), combined with VLMs, can learn to perform tasks zero-shot from new prompts, but they need to sample observations and actions from the dataset for computing rewards and for policy learning.
GenRL (Fig. 9c) needs to sample (sequences of) observations from the dataset, to infer the initial latent states for learning in imagination. Afterwards, rewards can be computed on the imagined latent sequences, enabling policy learning.
Data-free GenRL (Fig. 9d), samples the initial latent states internally by combining: (i) random samples of the latent space, (ii) randomly sampled embeddings, which are mapped to “actual embeddings" using the aligner, and turned into latent states, by the connector. Thus, policy learning requires no data sampling at all.
Appendix I Scaling to complex observations
Figure 10: Decoded language prompts in Minecraft
Generalist embodied agents should be able to scale to open-ended learning settings. Using GenRL, we explored this by training an agent in the Minecraft environment using a small dataset collected by a DreamerV3 agent [26]. Note that GenRL is also trained using similar experimental settings as DreamerV3, e.g. setting up the environment in the same way. The primary challenge we found was the model’s difficulty in reconstructing complex observations in this open-ended environment.
Reconstructing complex observations is a common issue with world models [10]. To overcome this limitation, while keeping the method unaltered, we attempted to scale up the number of parameters of the MFWM. Qualitative reconstruction results are presented in Figure 10. We observe that the agent is able to identify different biomes from language, even with the smaller size of the model. However, the reconstructions are significantly blurrier compared to the other environments we analyzed. When using a larger model, the reconstructions gain some details but the results still highlight the difficulty of the model in providing accurate targets from prompts.
While this might not be an issue for simple high-level tasks, e.g. ‘navigate to a beach’, inferring unclear targets might make it difficult to perform more precise actions, e.g. ‘attack a zombie’. Future research should aim to address this issue, for instance, by improving our simple GRU-based architecture, leveraging transformers or diffusion models to improve the quality of the representation [30, 3].
Appendix J Additional limitations
Pre-training data requirements. As we developed our framework, we observed that, in order to solve more complex tasks, the agent requires some expert data/demonstration of the complex behavior. We observed and analyzed this aspect in the experiments of Section 4.3. We believe this limitation is, to some extent, inevitable, as data-driven AI agents need to observe complex behaviors during training in order to be able to replicate them.
In this work, we used some exploration data, obtained using an exploration agent (Plan2Explore), and some task-specific data, collected using an expert RL agent (DreamerV3). Alternatively, one could investigate the adoption of a small set of demonstrations.
Precise manipulation skills. Adding additional tasks to the Kitchen environment, for testing multi-task generalization, showed to be harder than for the other domains. This is due to the difficulties of exploring meaningful behaviors in manipulation environments.
We are able to use GenRL to retrieve the four tasks present in the dataset and we are generally able to achieve similar tasks, such as “reach the microwave". From decoding the language prompts, we also see that GenRL can sensibly decode prompts such as “moving to the left" or “staying still", or more ’unusual’ prompts such as “swan pose" (where the robot arm would imitate the pose of a swan). While these prompts show that the system works well in this environment, none of these tasks would be particularly useful.
When we prompt the system with more interesting tasks that are out of the distribution, the agent generally struggles. For instance, if we ask to open the double door cabinet on the top left, the agent may reach for it but will not open it. The reason behind this is that it probably never opened that cabinet in the training dataset and thus it’s impossible to reproduce such behavior in an offline manner. The same issues are most likely present with all the baselines as well, as they all use the same dataset and foundation models for training.
For the locomotion domains, instead, the exploration data is varied enough that new tasks, outside of the tasks present in the training dataset, are enabled through the exploration data/
Appendix K Detailed results
@@ Appendix I Scaling to complex observations
Figure 11: Multi-task generalization detailed results. Results averaged over 10 seeds.
@@
Figure 12: Data-free RL detailed results. Results averaged over 10 seeds.
@@
Figure 13: Training data distribution detailed results. Results averaged over 10 seeds.